# GlossLM: A Massively Multilingual Corpus and Pretrained Model for Interlinear Glossed Text

### v2

@@ Abstract
Furthermore, we explore the task of automatically generating IGT in order to aid documentation projects. As many languages lack sufficient monolingual data, we pretrain a large multilingual model on our corpus. We demonstrate the utility of this model by finetuning it on monolingual corpora, outperforming SOTA models by up to 6.6%. We will make our pretrained model and dataset available through Hugging Face, as well as provide access through a web interface for use in language documentation efforts.
@@ 1 Introduction
Though IGT often follows a common glossing format, gloss conventions vary wildly. Furthermore, IGT data is rarely compiled into large, standardized corpora, often existing as scattered examples in research papers. To address this, we compile the largest corpus of digitized IGT from various existing sources, with over 450k examples in 1.8k languages (§3). We explore methods to normalize this data (§4)., standardizing over 80% of the grammatical glosses in the corpus to follow the UniMorph (Sylak-Glassman, 2016). We will release our corpora for future NLP, linguistics, and language documentation research.
@@
Futhermore, creating new IGT corpora is difficult and time-consuming, requiring documenters to perform linguistic analysis and extensive documentation simultaneously. Research has found that computational tools can help accelerate annotation and overcome this bottleneck Palmer et al. (2009) by predicting the gloss line of IGT given a transcription.
@@
We aim to overcome the monolingual data bottleneck by creating a multilingual pretrained glossing model that can be adapted to specific languages and gloss formats. We continually pretrain a model on our corpus, finding that pretrained multilingual model retains high performance across languages. We then finetune the pretrained model on monolingual data, including languages that do not appear in the pretraining corpus. We achieve a new SOTA performance on five out of seven languages, and demonstrate clear improvements for low-resource language settings over an equivalent finetuned model without our continual pretraining (§7).
@@ 3 GlossLM Corpus
We compile and clean the largest IGT dataset from a variety of sources and languages. In total, our dataset contains over 450k IGT instances (from 250k unique sentences) in 1.8k languages, collected from six different IGT corpora. All sources are publicly available under the CC BY 4.0 License, allowing free use and redistribution, and we have confirmed with the creators of each source that our usage is within the intended use. We make our artifacts available under the same license.
@@ 4.1 Motivation
As our data comes from a variety of sources, spanning many languages and documentation projects, there is a great amount of diversity of the morphological glosses used. This includes cases where several different labels are used to indicate the same feature (e.g. Sing, Sg, or S for singular), as well as formatting differences such as the usage of periods (e.g. 1Sg vs 1.Sg).
We explore the value and feasibility of normalizing glosses to a single standardized format. On one hand, normalizing glosses may make it easier to train models that utilize crosslingual information through shared gloss labels, but on the other hand, it can be difficult to select a single schema that preserves the original intent of all annotators.
We split gloss lines by period and count the number and frequency of unique grammatical gloss labels across our corpus (focusing on the all-caps functional glosses, not stems) and visualize the distribution in Figure 2.
@@ 4.2 Methodology
Several labels were ambiguous, corresponding to one of several UniMorph glosses depending on the language and annotator. For example, the label S (appearing 20,855 times) is used for singular, subject, and noun/sustantivo (at least) in our dataset. In order to map these, we would need to analyze on a sentence-by-sentence basis, which was not feasible; thus, we left these labels as-is.
@@
We use our mapping to normalize the dataset and will make the normalized version available in addition to the original.
@@ 4.3 Use in Future Research
Linguistics researchers will be able to use the dataset to easily search for phenomena across languages, particular with the normalized version of the dataset. For example, a linguist could find examples of sentences demonstrating the ergative/absolutive distinction. They could further refine this analysis by narrowing the results to related languages, using the glottocodes in our dataset.
@@ 7.2 Comparison with Finetuned ByT5
With enough data (starting with Nyangbo, 2100 examples), the two approaches achieve nearly identical performance. This is an unsurprising result, indicating that large amounts of high-quality monolingual data overshadow any benefits from crosslingual transfer. Furthermore, we note that benefits of multilingual gloss pretraining shown may not be unique to the T5 architecture—while we only experiment with ByT5, our pretraining strategy could be applied to other architectures.
@@ 7.3 Effect of Gloss Normalization
We repeat the same pretraining and finetuning process as before. When comparing the performance of the pretrained model before finetuning on in-domain languages, we find minimal differences compared to the results in §6.3. This is also the case after finetuning, though we found improvements in three of the four out-of-domain languages: Lezgi ($+\Delta 2.9$), Natugu ($+\Delta 6.1$), and Nyangbo ($+\Delta 2.5$), which (aside from Gitksan) are the three lowest resourced languages in our test set, exceeding current SOTA performance. From this, it appears that normalization may be most helpful when adapting the pretrained model to an unseen language with a low-to-moderate amount of training data.
@@ Appendix A Language Distribution
Figure 6: Counts per language. We only show languages with at least 2k samples present in the dataset. Arapaho (arap1274) is by far the most represented language in our data, followed by Uspanteko (uspa1245). Both languages are part of the SIGMORPHON Shared Task dataset.
Figure 7: Counts per language family in our dataset. We only show language families with at least 1k samples present in the dataset. Language isolates and languages without recorded families in Glottolog Hammarström et al. (2023) are categorized under “None”.
In Figure 6 and Figure 7, we visualize the number of examples per language and language family on a portion of our dataset.
@@ Appendix E In- vs out-of-vocabulary errors
Appendix F Example System Outputs
@@
We include examples of system outputs to illustrate the types of errors discussion in Table 10.

### v3

@@ Abstract
Furthermore, we explore the task of automatically generating IGT in order to aid documentation projects. As many languages lack sufficient monolingual data, we pretrain a large multilingual model on our corpus. We demonstrate the utility of this model by finetuning it on monolingual corpora, outperforming SOTA models by up to 6.6%. Our pretrained model and dataset are available on Hugging Face. [^1] https://huggingface.co/collections/lecslab/glosslm-66da150854209e910113dd87
@@ 1 Introduction
Though IGT often follows a common glossing format, gloss conventions vary wildly. Furthermore, IGT data is rarely compiled into large, standardized corpora, often existing as scattered examples in research papers. To address this, we compile the largest corpus of digitized IGT from various existing sources, with over 450k examples in 1.8k languages (§3). We explore methods to normalize this data (§4)., standardizing over 80% of the grammatical glosses in the corpus to follow the UniMorph schema (Sylak-Glassman, 2016). We are releasing our corpora for future NLP, linguistics, and language documentation research.
@@
The creation of new IGT corpora is often difficult and time-consuming, requiring documenters to perform linguistic analysis and extensive documentation simultaneously. Research has found that computational tools can help accelerate annotation and overcome this bottleneck Palmer et al. (2009) by predicting the gloss line of IGT given a transcription.
@@
We aim to overcome the monolingual data bottleneck by creating a multilingual pretrained glossing model that can be adapted to specific languages and gloss formats. We continually pretrain a model on our corpus, and find that the pretrained multilingual model retains high performance across languages. We then finetune the pretrained model on monolingual data, including languages that do not appear in the pretraining corpus. Our models achieve new SOTA performance on five out of seven languages, and demonstrate clear improvements for low-resource language settings over an equivalent finetuned model without our continual pretraining (§7).
@@ 3 GlossLM Corpus
While various publicly available sources of digitized IGT exist, the lack of unified data formatting and ease of access is a roadblock to using these resources effectively. To solve this problem, we compile and clean the largest IGT dataset from a variety of sources and languages. In total, our dataset contains over 450k IGT instances (from 250k unique sentences) in 1.8k languages, collected from six different IGT corpora. All sources are publicly available under the CC BY 4.0 License, allowing free use and redistribution, and we have confirmed with the creators of each source that our usage is within the intended use. We make our artifacts available under the same license.
@@ 4.1 Motivation
As our data comes from a variety of sources, spanning many languages and documentation projects, there is a great amount of diversity in the morphological glosses used. This includes cases where several different labels are used to indicate the same feature (e.g. Sing, Sg, or S for singular), as well as formatting differences such as the usage of periods (e.g. 1Sg vs 1.Sg).
We explore the feasibility and value of normalizing glosses to a single standardized format. On one hand, normalizing glosses may make it easier to train models that utilize crosslingual information through shared gloss labels, but on the other hand, it is difficult (perhaps impossible) to select a single schema that preserves the original intent of all annotators.
We split gloss lines by period and count the number and frequency of unique grammatical gloss labels across our corpus (focusing on the all-caps functional glosses, not stem translations) and visualize the distribution in Figure 2.
@@ 4.2 Methodology
Several labels were ambiguous, corresponding to one of several UniMorph glosses depending on the language and annotator. For example, the label S (appearing 20,855 times) is used for singular, subject, and noun/sustantivo (at least) in our dataset. In order to map these, we would need to analyze their meaning on a sentence-by-sentence basis, which was not feasible; thus, we left such ambiguous labels as-is.
@@
We use our mapping to normalize the dataset and make the normalized version available in addition to the original.
@@ 4.3 Use in Future Research
Linguistics researchers will be able to use the dataset to easily search for phenomena across languages, particularly with the normalized version of the dataset. For example, a linguist could find examples of sentences demonstrating the ergative/absolutive distinction. They could further refine this analysis by narrowing the results to a set of related languages, using the glottocodes in our dataset.
@@ 6.3 Performance of Pretrained Model
The languages evaluated here are well-represented in the pretraining corpus, from Tsez (3.7k unsegmented examples, about 3% of the total corpus) to Arapaho (39.1k examples, 21%). A natural question is whether the model retains good performance on a language which occurs very rarely in the pretraining corpus. We simulate this scenario by adding a small amount of data to the pretraining corpus for two unseen languages: 1000 examples in Nyangbo and 500 in Natugu (less than 1% of the total corpus). We evaluate on the unseen test split and observe 76.8% and 55.0% morpheme accuracy, respectively. These results indicate the model can still perform well on languages that are sparse in the pretraining corpus.
@@ 7.1 Comparison with Baselines
To illustrate common errors from the finetuned GlossLM moderls, we include examples of system outputs in Table 10. When inspecting outputs, we observe that there are sometimes inconsistencies in the IGT labels where multiple interchangeable glosses are used for the same morpheme. While we try to account for a portion of grammatical gloss variations (§4, §7.3), this is particularly an issue for lexical glosses (e.g. the Arapaho word ’eeneisih’i is glossed in the data as “how.X.things.are.named”, “how.s.o..is.named”, and “how.things.are.named”). We also find outputs that include lexical items present in the translation that are not included in the gold gloss, indicating that the model may rely too heavily on translations when predicting lexical glosses in certain cases.
@@ 7.2 Comparison with Finetuned ByT5
For Lezgi—which shows the greatest improvements from pretraining with the GlossLM corpus—a qualitative analysis of examples with the greatest morpheme error rate between finetuned ByT5 and GlossLM reveals that there are regular error patterns that are fixed with continual pretraining. For example, the finetuned ByT5 model often outputs AOC instead of AOR and OLB instead of ERG, whereas the finetuned GlossLM gets these correct. We include examples of these outputs in Table 11.
However, with enough data (starting with Nyangbo, 2100 examples) the two approaches achieve nearly identical performance. This is an unsurprising result, indicating that large amounts of high-quality monolingual data overshadow any benefits from crosslingual transfer. Furthermore, we note that benefits of multilingual gloss pretraining shown may not be unique to the T5 architecture—while we only experiment with ByT5, our pretraining strategy could be applied to other architectures.
@@ 7.3 Effect of Gloss Normalization
We repeat the same pretraining and finetuning process as before. When comparing the performance of the pretrained model before finetuning on in-domain languages, we find minimal differences compared to the results in §6.3.
Figure 6: Change in morpheme accuracy after normalizing glosses to the UniMorph schema and finetuning GlossLM. We observe small improvements for several languages, but worse performance in two cases.
We report the change in morpheme accuracy after normalizing in Figure 6. We observe mixed results. Some languages (Arapaho, Uspanteko, Gitksan) show worse or equivalent performance. Others (Tsez, Nyangbo, Natugu, and Lezgi) show small to moderate improvements, with Lezgi achieving the largest improvement of 3.0 percentage points. Thus, we found that normalization was most helpful when finetuning the pretrained model on unseen languages with a low-to-moderate amount of training data.
@@ 9 Conclusion
Acknowledgements
We would like to thank the three anonymous reviewers and meta-reviewer for their helpful feedback on this work. Parts of this work were supported by the National Science Foundation under Grant No. 2149404, “CAREER: From One Language to Another” and Grant No. 2211951, "From Acoustic Signal to Morphosyntactic Analysis in One End-to-End Neural System." Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Language Distribution
Figure 7 and Figure 8 display the number of examples per language and language family on a portion of our dataset.
@@ Appendix E In- vs out-of-vocabulary errors
Appendix F Example Outputs
We include example outputs to show the errors discussed in §7.1 and §7.2.
@@ Appendix F Example Outputs
Sample ID MER
lezg1247_71 transcription Акъадарда и пачагь ибуру ламрал , балкІандал акъадарда , ламрал акъадарда яда , цІайни ахъайда , им гьи хуьрей агъуз .
gold mount.ENT this king these-ERG donkey.on horse-SPSS mount.ENT donkey.on mount.ENT or fire.and released.ENT he this village.out.of down
ByT5 mount.ENT-ENT this king these-ERG donkey-OBL-SPSS horse-OBL-SPSS mount.ENT-ENT donkey-OBL-SPSS mount.ENT-ENT was fire-FOC 0.585
   put.ENT-ENT he this village-INESS-SPSS down.ENT-ENT
GlossLM mount.ENT this king these-ERG donkey.on horse-SPSS mount.ENT donkey.on mount.ENT or fire.and mount.ENT he this village.out.of down 0.068
lezg1247_49 transcription Хтана балкІан , « Гьаа », лагьана « гила чавай физ жеда » лагьана , « пачагьдин руш гъиз ».
gold return-AOR horse Yes say-AOR now help-INELAT go-DAT be-ENT say-AOR king-DIR-GEN girl will.bring-DAT
ByT5 coming-AOC horse Yes say-AOC now king-DIR-GEN girl be-ENT say-AOC king-DIR-GEN girl be-ENT 0.398
GlossLM return-AOR horse Yes say-AOR now help-INELAT go-DAT be-ENT say-AOR king-DIR-GEN girl will.bring-DAT 0
lezg1247_27 transcription АтІуз гана ибурун кьилерни .
gold cutting.IMC-DAT give-AOR these-ERG-GEN chapter-PL-FOC
ByT5 then give-AOR these-ERG-GEN head-PL-FOC 0.380
GlossLM cutting.IMC-DAT give-AOR these-ERG-GEN chapter-PL-FOC 0
lezg1247_0 transcription И гададини гъил вегьена са жуьт къачуда .
gold this boy-DIR-GEN-ERG hand threw-AOR one pair take-INESS
ByT5 this boy-DIR-GEN-ERG hand took-AOC one necklace take-ENT 0.327
GlossLM this boy-DIR-GEN-ERG hand threw-AOR one pair take-INESS 0
lezg1247_42 transcription И рушни пачагь хьана гила башламишда вичин пачагьвализ .
gold this girl-Q king happened-AOR now started.ENT-INESS himself-ERG-GEN reigned.ENT-ERG-DAT
ByT5 this girl-FOC king be-AOC now started.ENT-ENT himself-OBL-GEN reign.to-OBL-DAT 0.325
GlossLM this girl-Q king happened-AOR now started.ENT-INESS himself-ERG-GEN reigned.ENT-ERG-DAT 0
Table 11: Lezgi examples with the highest difference in MER between finetuned ByT5 and GlossLM outputs.